Grade 11General Physics

Two identical stell cubes of masses 50g and 1cm;collide head-on;face to face,with a speed of 10cm/each.Find the maximum compression of each.Young ‘s modulus for steel=2*10^11 N/m^2

Vidit
10 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
.Young ‘s modulus = Fl/Adl
50*10/1000*0.01/(0.01)^2*dl = 2*10^11 N/m^2
get the value of dl from here.that would be the maximum compression.